# An Overview of 7726 User Reports: Uncovering SMS Scams and Scammer Strategies

### v1

@@ IV-C1 Sender IDs
Table VII shows that more than 99% phone number scammers abuse for the Wrong Number, Hi Mum/Dad, and Government scams originate in the UK While more than 91% of phone numbers originate from the UK for Delivery and Banking scams, we notice phone numbers originating from the USA and the Philippines abused for Delivery scams and numbers from Japan abused for Banking scams. This is likely due to preferences of a few attackers. Scam campaigns impersonating only two particular UK banks account for almost all of the Japanense number (8.49% of the total banking scams). Similarly, we identify a campaign impersonating one Delivery brand with nearly identical text explaining most of the US numbers (and 3.09% of delivery scams) and two campaigns originating from the Philippines, one of which belongs to the ‘darcula’ iMessage and RCS smishing attacks [62].
@@
We find that the median lifetime of all phone numbers scammers abuse is 48 days (6.86 weeks). While 75.3% are active after 2 weeks, only 35.8% are active after 10 weeks. For Hi Mum/Dad scams, we find the median to be 4.14 weeks, i.e., 29 days, whereas previous research found the median lifetime to be 14 days [30]. We attribute this difference to the fact that previous work only studies scam messages detected by existing XDR filters, missing those that permeate through their detection systems. Instead, our work investigates phone numbers reported by users. While we discover that the median lifetime of phone numbers abused to conduct delivery and banking scams is the same as the Hi Mum/Dad scams, the median lifetime for telecom is 9 weeks and over 9 weeks for wrong number scams. This indicates scammers being able to evade mobile network operator’s XDR systems.
@@ IV-C2 Mobile Network Operators
UK mobile network operators that provide physical SIM cards are more abused than Virtual mobile network operators. This could be because scammers broadcast thousands of scam texts using a SIM box/farm [63]. A SIM box is an SMS gateway device connected to a computer with multiple SIM slots. MNO 1 is the most preferred mobile network operator, followed by MNO 2, 3, and 4. While MNO 1, 2, and 3 support GSM technology that works with SIM boxes, MNO 4 does not support GSM. The desirability of mobile network operators by scammers could also depend on the ease of availability to procure the SIM cards, network availability where the scammers are based, or the XDRs that scammers evade to detect these scam texts.
@@
Next, we explore the popularity of mobile network operators based on scam types. While MNO 1 is least abused when sending Telecom scams, MNO 2, 3, and 4 are least abused when it comes to Hi Mum/Dad and Government scams. This could mean that MNO 1 is better at detecting Telecom scams, whereas MNO 2, 3, and 4 are better at blocking Hi Mum/Dad and Government scams. While most Virtual numbers are primarily abused to send Wrong Number scams, MVNOs are abused for Telecom, Hi Mum/Dad, and Wrong Number scams. Virtual numbers provide scammers the advantage of sending scam texts from outside the UK. However, with MVNOs, even though they use the network of a physical MNO, the text messages go through their own XDR systems (if implemented). Looking into MNO 10, the only non-UK mobile network operator in Table VIII, we find similar text messages targeting one Delivery company. This indicates that a single threat actor abused 3.09% of all sender IDs that belong to the USA. For Hi Mum/Dad scams, MNO 1 is significantly more abused (86.4%) than others [30] showcasing that scammers abuse different types of mobile network operators for various scams and some more than others.
@@ IV-D Domain Analysis
URL Shorteners. We identify a significant amount of URLs in the text message reports that are shortened URLs, inline with prior research [32]. Table IX presents the number of unique URLs belonging to the 10 most popular URL shorteners abused to conduct scams, with the two most common services being bit.ly and t.ly. Scammers abuse URL shortening services to evade detection from XDRs and antivirus vendors. Shortened URLs also can be created for free, fit in an SMS character limit, and redirect to the malicious URL only when a user clicks.
@@
Registrars. We find that NameSilo (26.3%) is the most abused registrar for SMS scam URLs, followed by Hosting Concepts B.V. (13.2%). Previous research on newly registered phishing domains also found NameSilo as the most abused registrar [18]. On the other hand, a smishing research using US smish reports identified NameCheap [33] as the most abused registrar. This highlights differences in our collected data as well as the impact of our spam/scam distinction. While domains in delivery and telecom scams primarily abuse NameSilo, MarkMonitor is the most abused registrar for banking and government scams. This highlights that different scams abuse different registrars, likely reflecting preferences of different groups of threat actors.
@@
We find that ‘.com’ remains the most abused TLD, followed by ‘.top’ and ‘.co.uk.’ This is in line with recent research [18, 64, 65, 66]. Investigating individually by scam types – ‘.com’ remains the most abused for delivery, banking, telecom, government, and wrong number scams. After ‘.com,’ we uncover that the delivery scams abuse .top followed by .xyz. On the other hand, we find that government scams abuse ‘co.uk’ and ‘.uk’ TLDs making them more convincing for potential victims. For example, arrange-test-kit[.]co[.]uk is used to impersonate a health service text. While ‘.buzz’ is the second most abused TLD by scammers for telecom scams, they abuse ‘web.app’ for banking scams. For example, scammers set up attempted-logon[.]web[.]app to lure victims into clicking on the malicious link.
@@
Autonomous Systems (ASes). We identify 56,092 unique IP addresses that the identified scam domains abused. We find that 7,110 (12.7%) unique IP addresses belong to Cloudflare, a proxy service which hides the server IP address. Previous research also found most domains in their dataset abusing Cloudflare [32, 18]. Table XI presents the five most abused Autonomous Systems (ASes) to host the identified scam URLs, excluding Cloudflare. It also shows the geographical location of the IP addresses abused to host scam domains.
@@
Domain Randomness. We evaluate the randomness of second-level domains (SLDs) abused in smishing attacks. We find that 60.3% SLDs contain a (nontrivial) dictionary word, indicating use of meaningful words. This may imply the use of words such as ‘cancel,’ ‘track,’ and ‘ship’ along with brand names (including typosquatted brand names), which would be consistent with common impersonation tactics inline with previous research [67, 18, 68, 69, 70]. While the mean Shannon entropy is 2.88, suggesting moderate randomness, the average vowel-to-consonant ratio is 0.64, higher than expected for randomly generated strings. This indicates a bias toward pronounceable patterns. We select a random sample of 100 domain names and find 67% to be brand typos (e.g., evrlgb-couriers, verify-myapplepay, hsbc-cancel-payment, icloud-uk), also known as combosquatting [70]. The rest contain arbitrary character sequences (e.g., dcfcy, onapuw, azirew, acozir).
Domain Lifetime. We find the lifetime of 5,854 unique domains (excluding URL shorteners), with a median of 118 days. We use survival analysis to investigate lifetime of all identified scam domains with overall survival probability (Fig. 7). The dotted lines in the plot are the 95% confidence interval. While 83.3% of the domains are active after 10 days, 65.8% are active after 100 days. This demonstrates that longer lived domains live longer. Previous work found that more than 40% of the domains from the US Federal Trade Commission (FTC) user reports were active for over 100 days [36]. On the contrary, Nahapetyan et al. observe an average lifetime of 12.241 days with a median of 0.53 hours [32]. Others identify the average lifetime of scam domains as 59.4 days [71].
@@
Particularly, the identified domains last so long as scammers register and re-register domain names, continuing to abuse them over time [72]. For example, we see uk-delivery[.]com was first seen on April 20, 2014, and has been likely re-registered multiple times since then.
@@ IV-E SMS Text Analysis
Scam Lures. Scammers design text messages to deceive victims, asking them to take action. We work to understand this deception by classifying the scams by lure using the typology identified by Stajano and Wilson [48]. Table XII presents the scams we classify by lures. Scammers use the time/urgency lure in all scam types, forcing users to make an impulsive decision. For example, using words like ‘now,’ ‘urgent,’ or ‘immediately.’ We find that scammers lure victims into Hi Mum/Dad scams by also employing distraction and kindness, similar to prior findings [30]. For example, scammers pretend to be a victim’s child, provide a random reason for reaching out from a different number and request immediate help. Similarly, for Wrong Number scams, scammers provide unrelated details and asking a random question or sounding like a friend to deceive potential victims into replying, which can turn into a pig butchering scam [73].
@@
TABLE XII: Scam types categorized by lures (adapted from Stajano and Wilson [48]).
Scammers use need and greed to lure victims into Government (12.6%), Telecom (13.7%) and Banking (11.7%) scams by offering tax refunds, points, cashback, or mentioning a suspicious payment, tempting the user to take the requested action in the text. Government (67%), Telecom (47.1%), Banking (44.2%), and Delivery (37.8%) scams also invoke authority, making it more convincing by impersonating government and private entities such as HMRC, EE, Barclays, or EVRI. This indicates that scammers do not just randomly draft scam texts but take advantage of various lures to deceive potential victims into taking an action. Depending on the type of scam, scammers use one or more lures in the text of the message. Unsurprisingly, we do not find SMS scammers using herd and dishonesty lures despite these commonly used in cryptocurrency scam ads [74, 75].
Gunning Fog Index. Scammers design the text of the scam messages depending on their target victims. The text message’s readability is essential as not all potential victims can understand complex sentences in scam text messages. To target a broader range of individuals, one would expect the scam text messages to be simple. Towards this end, we use the Gunning Fog Index (GFI), designed to compute the years of education required to understand a given passage/text [47]. We find that the mean of the Gunning Fog Index for conversational scam text messages is lower than others. While Hi Mum/Dad and Wrong number scams have a Gunning Fog Index of 4.6 and 3.9, respectively, Delivery/Parcel, Banking, Telecom, and Government impersonation scams have the Gunning Fog Index above 5.6, the maximum being 8.3 for Delivery/Parcel scams (Fig 8 in Appendix shows the GFI for all scam types). The Gunning Fog Index does not significantly differ between scam types. This indicates that scammers design straightforward scam texts that anyone with a maximum of 8th-grade education can read and understand.
@@ V Discussion
Differentiating fraudulent from non-fraudulent messages is crucial in an increasingly challenging landscape. Our analysis reveals that over 40% of text messages are scams, observing on the other end a significant number of spam reports. However, we see that messages with fraudulent intent show different patterns than those without. Quite a number of our findings vary with intent like the likelihood of using a phone number vs. other sender IDs or the likelihood of using a URL. Previous research fails to differentiate between the scam and spam text messages, limiting its effectiveness in combating smishing [36, 31, 33, 32]. This distinction is of heightened importance when factors like the infrastructure or text patterns used are key ones that feed into XDRs at mobile network operators. When these factors differ, stakeholders require to adopt distinct strategies to address fraudulent messages effectively using user reports [15]. We call upon other researchers to also make this vital distinction in order to positively protect users from harm.
@@
Scammers shift toward non-URL-based lures, which bypass mobile network operators’s XDR systems. With the uptake in SMS scams, we see “Wrong Number” and other conversational scams dominating the submissions — here we identify six new scam types above narrower prior work [19, 30]. These new types of conversational scams bypass mobile network operators’ XDR systems, which are more effective at detecting URL-based threats. This indicates a critical gap in current detection mechanisms, as most reported scams are texts without URLs, and shows the need to enhance algorithms for identifying such messages. This feature also highlights a key distinction from email phishing, posing challenges to the adoption of effective defense mechanisms [76, 77]. The shift in SMS scams echoes an evasion strategy similar to technical support scams, which have advanced from URL-based scareware URLs [78] to call-based attacks via malvertising [79], bypassing conventional URL-focused defenses.
@@
We note that Telecom regulators maintain a list of phone numbers (aka do not originate or DNO) from organizations like banks to combat call spoofing [80]. We suggest that mobile network operators query phone numbers in the text of an SMS that posits to be an organization against the DNO list as an effective legitimacy checking mechanism, in the direction of Domain-based Message Authentication, Reporting & Conformance (DMARC) used in mail servers [81] and telephony blocklists [82].
@@
Exploiting online encrypted messaging platforms to send scam texts. We also notice scammers use email addresses for delivery scams indicating RCS/iMessage scams. Rich communication service (RCS) and iMessage allow users to send encrypted SMS over the internet instead of SMS over mobile network operator. In addition to emails, we identify a scam campaign with over 44 deliver scam texts known to send RCS/iMessage scam texts called ‘darcula’ — a phishing-as-a-service platform [62, 83]. As mobile network operators improve their XDRs, scammers are turning to similar and cheaper alternatives like iMessage and RCS to bypass detection. The shift in smishing from traditional SMS to encrypted messaging platforms mirrors how traditional email-based phishing [84] has evolved towards cryptocurrency platforms [85, 86]. As RCS/iMessage texts directly bypass the mobile network operator’s XDRs, we suggest services like Apple and Google collaborate with mobile network operators to identify and proactively block scams, enabling a more coordinated and comprehensive mitigation strategy.
@@
Limitations. We receive the user reports from only one mobile network operator. The unavailability of SMS scam data is a general limitation here. We make significant progress on this limitation for one mobile network operator by devising a processing pipeline that can characterize popular, known scam types and identify novel ones. As a key contribution, we see that the lures scammers use to deceive victims into taking an action differs based on the scam type. Understanding the lures used could assist interveners to create more effective warnings and better user education about ongoing scams. Interveners include private companies like the Students Loans Company [87] and governments [88].
@@
While we identify spam and scams, we also flag 23.26% text reports as ‘Unknown.’ Even though the URLs in the reports are not flagged by an AV vendor, previous research has shown that more than 94% of the blacklisted domain names do not appear in public blacklists for several weeks or even months after they are first reported in abuse complaints [36]. There is room for improving our text characterization and scam type identification. Some of the misclassifications we observe reveal fraudulent messages mimicking legitimate messages, with some being as simple as ‘Hi’ — requiring more sophisticated detection mechanisms.
@@
However, the scope of our paper is not to devise such detection mechanisms, but to offer a first look into SMS scams and scammer strategies through the lens of a large base of recent user reports. Our evaluation metrics (88% for spam messages, 89% for known scam types, 80% new types) demonstrate high standards to answer our research questions, especially considering the scale and real-world conditions of our evaluation. While a deeper understanding of unknowns and other new types of scams is the scope of our future work, we reveal six emerging scam types (that group together) (§IV-B); finer breakdowns will be a valuable future direction. A challenge to address in this direction is to adopt novel detection mechanisms [89] to identify and classify new trends.

### v2

@@ IV-C1 Sender IDs
Table VII shows that more than 99% phone number scammers abuse for the Wrong Number, Hi Mum/Dad, and Government scams originate in the UK While more than 91% of phone numbers originate from the UK for Delivery and Banking scams, we notice phone numbers originating from the USA and the Philippines abused for Delivery scams and numbers from Japan abused for Banking scams. This is likely due to preferences of a few attackers. Scam campaigns impersonating only two particular UK banks account for almost all of the Japanense number (8.49% of the total banking scams). Similarly, we identify a campaign impersonating one Delivery brand with nearly identical text explaining most of the US numbers (and 3.09% of delivery scams) and two campaigns originating from the Philippines, one of which belongs to the ‘darcula’ iMessage and RCS smishing attacks [darcula_netcraft].
@@
We find that the median lifetime of all phone numbers scammers abuse is 48 days (6.86 weeks). While 75.3% are active after 2 weeks, only 35.8% are active after 10 weeks. For Hi Mum/Dad scams, we find the median to be 4.14 weeks, i.e., 29 days, whereas previous research found the median lifetime to be 14 days [agarwal_usenix25]. We attribute this difference to the fact that previous work only studies scam messages detected by existing XDR filters, missing those that permeate through their detection systems. Instead, our work investigates phone numbers reported by users. While we discover that the median lifetime of phone numbers abused to conduct delivery and banking scams is the same as the Hi Mum/Dad scams, the median lifetime for telecom is 9 weeks and over 9 weeks for wrong number scams. This indicates scammers being able to evade mobile network operator’s XDR systems.
@@ IV-C2 Mobile Network Operators
UK mobile network operators that provide physical SIM cards are more abused than Virtual mobile network operators. This could be because scammers broadcast thousands of scam texts using a SIM box/farm [dcpcu_simfarm]. A SIM box is an SMS gateway device connected to a computer with multiple SIM slots. MNO 1 is the most preferred mobile network operator, followed by MNO 2, 3, and 4. While MNO 1, 2, and 3 support GSM technology that works with SIM boxes, MNO 4 does not support GSM. The desirability of mobile network operators by scammers could also depend on the ease of availability to procure the SIM cards, network availability where the scammers are based, or the XDRs that scammers evade to detect these scam texts.
@@
Next, we explore the popularity of mobile network operators based on scam types. While MNO 1 is least abused when sending Telecom scams, MNO 2, 3, and 4 are least abused when it comes to Hi Mum/Dad and Government scams. This could mean that MNO 1 is better at detecting Telecom scams, whereas MNO 2, 3, and 4 are better at blocking Hi Mum/Dad and Government scams. While most Virtual numbers are primarily abused to send Wrong Number scams, MVNOs are abused for Telecom, Hi Mum/Dad, and Wrong Number scams. Virtual numbers provide scammers the advantage of sending scam texts from outside the UK. However, with MVNOs, even though they use the network of a physical MNO, the text messages go through their own XDR systems (if implemented). Looking into MNO 10, the only non-UK mobile network operator in Table VIII, we find similar text messages targeting one Delivery company. This indicates that a single threat actor abused 3.09% of all sender IDs that belong to the USA. For Hi Mum/Dad scams, MNO 1 is significantly more abused (86.4%) than others [agarwal_usenix25] showcasing that scammers abuse different types of mobile network operators for various scams and some more than others.
@@ IV-D Domain Analysis
URL Shorteners. We identify a significant amount of URLs in the text message reports that are shortened URLs, inline with prior research [nahapetyan2024sms]. Table IX presents the number of unique URLs belonging to the 10 most popular URL shorteners abused to conduct scams, with the two most common services being bit.ly and t.ly. Scammers abuse URL shortening services to evade detection from XDRs and antivirus vendors. Shortened URLs also can be created for free, fit in an SMS character limit, and redirect to the malicious URL only when a user clicks.
@@
Registrars. We find that NameSilo (26.3%) is the most abused registrar for SMS scam URLs, followed by Hosting Concepts B.V. (13.2%). Previous research on newly registered phishing domains also found NameSilo as the most abused registrar [agarwal25:weis]. On the other hand, a smishing research using US smish reports identified NameCheap [smishtank_timko24] as the most abused registrar. This highlights differences in our collected data as well as the impact of our spam/scam distinction. While domains in delivery and telecom scams primarily abuse NameSilo, MarkMonitor is the most abused registrar for banking and government scams. This highlights that different scams abuse different registrars, likely reflecting preferences of different groups of threat actors.
@@
We find that ‘.com’ remains the most abused TLD, followed by ‘.top’ and ‘.co.uk.’ This is in line with recent research [agarwal25:weis, korczynski2017reputation, domainz_sp16, kumar2021diversity]. Investigating individually by scam types – ‘.com’ remains the most abused for delivery, banking, telecom, government, and wrong number scams. After ‘.com,’ we uncover that the delivery scams abuse .top followed by .xyz. On the other hand, we find that government scams abuse ‘co.uk’ and ‘.uk’ TLDs making them more convincing for potential victims. For example, arrange-test-kit[.]co[.]uk is used to impersonate a health service text. While ‘.buzz’ is the second most abused TLD by scammers for telecom scams, they abuse ‘web.app’ for banking scams. For example, scammers set up attempted-logon[.]web[.]app to lure victims into clicking on the malicious link.
@@
Autonomous Systems (ASes). We identify 56,092 unique IP addresses that the identified scam domains abused. We find that 7,110 (12.7%) unique IP addresses belong to Cloudflare, a proxy service which hides the server IP address. Previous research also found most domains in their dataset abusing Cloudflare [nahapetyan2024sms, agarwal25:weis]. Table XI presents the five most abused Autonomous Systems (ASes) to host the identified scam URLs, excluding Cloudflare. It also shows the geographical location of the IP addresses abused to host scam domains.
@@
Domain Randomness. We evaluate the randomness of second-level domains (SLDs) abused in smishing attacks. We find that 60.3% SLDs contain a (nontrivial) dictionary word, indicating use of meaningful words. This may imply the use of words such as ‘cancel,’ ‘track,’ and ‘ship’ along with brand names (including typosquatted brand names), which would be consistent with common impersonation tactics inline with previous research [agten2015seven, agarwal25:weis, moore2010measuring, szurdi2014long, kintis2017hiding]. While the mean Shannon entropy is 2.88, suggesting moderate randomness, the average vowel-to-consonant ratio is 0.64, higher than expected for randomly generated strings. This indicates a bias toward pronounceable patterns. We select a random sample of 100 domain names and find 67% to be brand typos (e.g., evrlgb-couriers, verify-myapplepay, hsbc-cancel-payment, icloud-uk), also known as combosquatting [kintis2017hiding]. The rest contain arbitrary character sequences (e.g., dcfcy, onapuw, azirew, acozir).
Domain Lifetime. We find the lifetime of 5,854 unique domains (excluding URL shorteners), with a median of 118 days. We use survival analysis to investigate lifetime of all identified scam domains with overall survival probability (Fig. 7). The dotted lines in the plot are the 95% confidence interval. While 83.3% of the domains are active after 10 days, 65.8% are active after 100 days. This demonstrates that longer lived domains live longer. Previous work found that more than 40% of the domains from the US Federal Trade Commission (FTC) user reports were active for over 100 days [srinivasan2016understanding]. On the contrary, Nahapetyan et al. observe an average lifetime of 12.241 days with a median of 0.53 hours [nahapetyan2024sms]. Others identify the average lifetime of scam domains as 59.4 days [kotzias2025ctrl+].
@@
Particularly, the identified domains last so long as scammers register and re-register domain names, continuing to abuse them over time [hao_imc13]. For example, we see uk-delivery[.]com was first seen on April 20, 2014, and has been likely re-registered multiple times since then.
@@ IV-E SMS Text Analysis
Scam Lures. Scammers design text messages to deceive victims, asking them to take action. We work to understand this deception by classifying the scams by lure using the typology identified by Stajano and Wilson [StajanoFrank2011_principles]. Table XII presents the scams we classify by lures. Scammers use the time/urgency lure in all scam types, forcing users to make an impulsive decision. For example, using words like ‘now,’ ‘urgent,’ or ‘immediately.’ We find that scammers lure victims into Hi Mum/Dad scams by also employing distraction and kindness, similar to prior findings [agarwal_usenix25]. For example, scammers pretend to be a victim’s child, provide a random reason for reaching out from a different number and request immediate help. Similarly, for Wrong Number scams, scammers provide unrelated details and asking a random question or sounding like a friend to deceive potential victims into replying, which can turn into a pig butchering scam [ordekian:ecrime24].
@@
TABLE XII: Scam types categorized by lures (adapted from Stajano and Wilson [StajanoFrank2011_principles]).
Scammers use need and greed to lure victims into Government (12.6%), Telecom (13.7%) and Banking (11.7%) scams by offering tax refunds, points, cashback, or mentioning a suspicious payment, tempting the user to take the requested action in the text. Government (67%), Telecom (47.1%), Banking (44.2%), and Delivery (37.8%) scams also invoke authority, making it more convincing by impersonating government and private entities such as HMRC, EE, Barclays, or EVRI. This indicates that scammers do not just randomly draft scam texts but take advantage of various lures to deceive potential victims into taking an action. Depending on the type of scam, scammers use one or more lures in the text of the message. Unsurprisingly, we do not find SMS scammers using herd and dishonesty lures despite these commonly used in cryptocurrency scam ads [gilberto_ecrime22, agarwal2023_fc].
Gunning Fog Index. Scammers design the text of the scam messages depending on their target victims. The text message’s readability is essential as not all potential victims can understand complex sentences in scam text messages. To target a broader range of individuals, one would expect the scam text messages to be simple. Towards this end, we use the Gunning Fog Index (GFI), designed to compute the years of education required to understand a given passage/text [gunning1969fog]. We find that the mean of the Gunning Fog Index for conversational scam text messages is lower than others. While Hi Mum/Dad and Wrong number scams have a Gunning Fog Index of 4.6 and 3.9, respectively, Delivery/Parcel, Banking, Telecom, and Government impersonation scams have the Gunning Fog Index above 5.6, the maximum being 8.3 for Delivery/Parcel scams (Fig 8 in Appendix shows the GFI for all scam types). The Gunning Fog Index does not significantly differ between scam types. This indicates that scammers design straightforward scam texts that anyone with a maximum of 8th-grade education can read and understand.
@@ V Discussion
Differentiating fraudulent from non-fraudulent messages is crucial in an increasingly challenging landscape. Our analysis reveals that over 40% of text messages are scams, observing on the other end a significant number of spam reports. However, we see that messages with fraudulent intent show different patterns than those without. Quite a number of our findings vary with intent like the likelihood of using a phone number vs. other sender IDs or the likelihood of using a URL. Previous research fails to differentiate between the scam and spam text messages, limiting its effectiveness in combating smishing [srinivasan2016understanding, tang:ccs22, smishtank_timko24, nahapetyan2024sms]. This distinction is of heightened importance when factors like the infrastructure or text patterns used are key ones that feed into XDRs at mobile network operators. When these factors differ, stakeholders require to adopt distinct strategies to address fraudulent messages effectively using user reports [spamshield_mavenir]. We call upon other researchers to also make this vital distinction in order to positively protect users from harm.
@@
Scammers shift toward non-URL-based lures, which bypass mobile network operators’s XDR systems. With the uptake in SMS scams, we see “Wrong Number” and other conversational scams dominating the submissions — here we identify six new scam types above narrower prior work [agarwal_imc24, agarwal_usenix25]. These new types of conversational scams bypass mobile network operators’ XDR systems, which are more effective at detecting URL-based threats. This indicates a critical gap in current detection mechanisms, as most reported scams are texts without URLs, and shows the need to enhance algorithms for identifying such messages. This feature also highlights a key distinction from email phishing, posing challenges to the adoption of effective defense mechanisms [oset:ecrime18, sahingoz2019machine]. The shift in SMS scams echoes an evasion strategy similar to technical support scams, which have advanced from URL-based scareware URLs [stone2012underground] to call-based attacks via malvertising [miramirkhani2016dial], bypassing conventional URL-focused defenses.
@@
We note that Telecom regulators maintain a list of phone numbers (aka do not originate or DNO) from organizations like banks to combat call spoofing [ofcom_dno]. We suggest that mobile network operators query phone numbers in the text of an SMS that posits to be an organization against the DNO list as an effective legitimacy checking mechanism, in the direction of Domain-based Message Authentication, Reporting & Conformance (DMARC) used in mail servers [kucherawy2015domain] and telephony blocklists [pandit2018towards].
@@
Exploiting online encrypted messaging platforms to send scam texts. We also notice scammers use email addresses for delivery scams indicating RCS/iMessage scams. Rich communication service (RCS) and iMessage allow users to send encrypted SMS over the internet instead of SMS over mobile network operator. In addition to emails, we identify a scam campaign with over 44 deliver scam texts known to send RCS/iMessage scam texts called ‘darcula’ — a phishing-as-a-service platform [darcula_netcraft, akyazi2021measuring]. As mobile network operators improve their XDRs, scammers are turning to similar and cheaper alternatives like iMessage and RCS to bypass detection. The shift in smishing from traditional SMS to encrypted messaging platforms mirrors how traditional email-based phishing [moore2007empirical] has evolved towards cryptocurrency platforms [chendissecting, chen_cryptophish]. As RCS/iMessage texts directly bypass the mobile network operator’s XDRs, we suggest services like Apple and Google collaborate with mobile network operators to identify and proactively block scams, enabling a more coordinated and comprehensive mitigation strategy.
@@
Limitations. We receive the user reports from only one mobile network operator. The unavailability of SMS scam data is a general limitation here. We make significant progress on this limitation for one mobile network operator by devising a processing pipeline that can characterize popular, known scam types and identify novel ones. As a key contribution, we see that the lures scammers use to deceive victims into taking an action differs based on the scam type. Understanding the lures used could assist interveners to create more effective warnings and better user education about ongoing scams. Interveners include private companies like the Students Loans Company [student_smishing] and governments [ftc_howto].
@@
While we identify spam and scams, we also flag 23.26% text reports as ‘Unknown.’ Even though the URLs in the reports are not flagged by an AV vendor, previous research has shown that more than 94% of the blacklisted domain names do not appear in public blacklists for several weeks or even months after they are first reported in abuse complaints [srinivasan2016understanding]. There is room for improving our text characterization and scam type identification. Some of the misclassifications we observe reveal fraudulent messages mimicking legitimate messages, with some being as simple as ‘Hi’ — requiring more sophisticated detection mechanisms.
@@
However, the scope of our paper is not to devise such detection mechanisms, but to offer a first look into SMS scams and scammer strategies through the lens of a large base of recent user reports. Our evaluation metrics (88% for spam messages, 89% for known scam types, 80% new types) demonstrate high standards to answer our research questions, especially considering the scale and real-world conditions of our evaluation. While a deeper understanding of unknowns and other new types of scams is the scope of our future work, we reveal six emerging scam types (that group together) (§IV-B); finer breakdowns will be a valuable future direction. A challenge to address in this direction is to adopt novel detection mechanisms [pimentel2014review] to identify and classify new trends.